OpenStudy (anonymous):

The MKS units of momentum, mass, and speed are, respectively, newton seconds, kilograms, and meters per second. Newtons are equal to kilogram meters per square seconds. Which of the following equations is incorrect by unit analysis? 1/m = v/p v = p/m p = mv m = p/v m = pv

OpenStudy (shamim):

For ur first option v/p=v/mv=1/m Right?

OpenStudy (shamim):

Nd it is equal to left hand side of ur first choice

OpenStudy (shamim):

So first choice is true

OpenStudy (shamim):

Now wt do u think abt ur second choice?